OpenStudy (anonymous):

V = 100 sin (200πt + π/4) the time to reach the greatest voltage rate of change

OpenStudy (anonymous):

That's cool. Now where's the Q?

OpenStudy (anonymous):

when is Sin theta maximum? for what angle?

OpenStudy (anonymous):

Oh, voltage rate of change...not voltage...then u ll need to differentiate first

OpenStudy (anonymous):

@telltoamit, I suspect your suspicions on the question is correct.

OpenStudy (anonymous):

question are correct.*

OpenStudy (anonymous):

yea, u ll jus get a Cos term

OpenStudy (anonymous):

is this correct dv/dt = (2 x 10^4 π) cos (200 π t + π/4)

OpenStudy (anonymous):

absolutely

OpenStudy (anonymous):

@tomtom, please write your questions as a question in the English language next time. It makes you more easily understood.

OpenStudy (anonymous):

what next after differentiation

OpenStudy (anonymous):

for what angle is Cos theta maximum?

OpenStudy (anonymous):

45?

OpenStudy (anonymous):

Nah. 0. Cos zero = 1

OpenStudy (anonymous):

is this my answer dv/dt = (2 x 104 π) cos (200 π t + π/4) maximum occurs when dv/dt = 0 and d2v/dt2 is negative d2v/dt2 = -(4 x 106 π2) sin (200 π t + π/4) cos (200 π t + π/4) = 0 when t = 0 and when t = 0 d2v/dt2 is negative when t = 0, v = 100 as sin (200 π + π/t) = 1 dv/dt is a maximum when cos (200 π t + π/4) is either 1 or -1 which occurs when 200 π t + π/4 = 0 or π radians t is either = -1/800 or 1/800 seconds t = -0.00125 or 0.00125 seconds

OpenStudy (anonymous):

No. U hv to find when is dv/dt maximum and not when v is maximum.

OpenStudy (anonymous):

maximum occurs when dv/dt = 0 so 0 = (2 x 10^4 π) cos (200 π t + π/4)

OpenStudy (anonymous):

No, the greatest voltage rate of change maximum for dv/dt occurs when d2v/dt2 = 0

OpenStudy (anonymous):

d2v/dt2 = (4 x 10^6 π2) sin (200 π t + π/4) = 0